# Multi-step Problem Solving Through a Verifier: An Empirical Analysis on Model-induced Process Supervision

### v1

@@ Abstract
Large language models often struggle to generate error-free solutions in complex problem-solving scenarios. To address this, recent advancements have adopted a reasoner-verifier framework, where a verifier model is trained in a supervised fashion, to evaluate intermediate solution steps generated by a reasoning model. However, this “process supervision” approach requires obtaining ground truth annotations on intermediate solution steps to train the verifier model, which is resource-intensive and expensive. In this paper, we introduce Model-induced Process Supervision (MiPS), a novel method for automating data curation. MiPS annotates an intermediate step by sampling completions of this solution through the reasoning model, and obtaining an accuracy defined as the proportion of correct completions. Errors in the reasoner would cause MiPS to underestimate the accuracy of intermediate steps, therefore, we suggest and empirically show that verification focusing on high predicted values of the verifier shall be preferred over that of low predicted values, contrary to prior work. Our approach significantly improves the performance of PaLM 2 on math and coding tasks (accuracy +0.67% on GSM8K, +4.16% on MATH, +0.92% on MBPP compared with an output verifier). Additionally, our study demonstrates that the verifier exhibits strong generalization ability across different reasoning models.
@@ 1 Introduction
Multi-step problem solving (e.g., math problems and coding challenges) showcases the capabilities of machine intelligence. While researchers have shown that model- and data-upscaling still hold powerful for large language models (LLMs) on multi-step problem solving Achiam et al. (2023); Touvron et al. (2023); Team Gemini et al. (2023); Huang et al. (2022); Azerbayev et al. (2023); Luo et al. (2023a); Yu et al. (2023a), even the state-of-the-art LLMs still produce easily observable mistakes. Furthermore, standard fine-tuning on the training dataset or model generated solutions directly does not yield consistent and significant improvements Luo et al. (2023a); Yu et al. (2023a); Ni et al. (2022). In Tab. 2, we show the test set performance of PaLM 2-S on a math problem dataset, GSM8K Cobbe et al. (2021), before and after task-specific finetuning. It is observed that such fine-tuning does not exploit the model’s full potential, as the performance falls much behind using a standard self-consistency approach Wang et al. (2022) and especially when considering that the model very likely generates a correct-answer solution to the problem.
Self-consistency can be seen as an inference-time technique where the goal is to pick one model-generated solution among many. This broadly falls under a reasoner-verifier paradigm (Fig. 2). In particular, the reasoner model generates solution(s) given a problem to solve and the verifier model makes predictions to the solutions, selecting the solution with the highest predicted value.
Training a verifier in a supervised fashion has demonstrated strong performance in both coding and math language problems. Cobbe et al. (2021) showed that by simply gathering correct and incorrect solutions to train a binary classification model and using such model to pick the highest confidence solution generated by the reasoner during inference time, the accuracy can be improved significantly. More recent studies suggest that verifying on intermediate steps could offer better guidance than training solely on the whole solutions Li et al. (2022); Uesato et al. (2022); Paul et al. (2023); Lightman et al. (2023); Feng et al. (2023); Yu et al. (2023b); Liu et al. (2023a); Wang et al. (2023). As such, verifiers trained (and applied) on intermediate steps are called process supervised verifiers (PSV), whereas those trained on whole solutions are called output supervised verifiers (OSV). In prior work, process supervision data is either obtained by an ad-hoc algorithm Li et al. (2022); Paul et al. (2023), or through expensive human annotations Uesato et al. (2022); Lightman et al. (2023), lacking an automatic and generic way of constructing of annotations of intermediate solutions.
@@
Figure 1: An illustration of the reasoner-verifier paradigm. The verifier predicts scores for the solutions generated by the reasoner, and selectes the solution with the highest score.
Method Acc.
Base, FT on Train (2K) 61.6
FT on Train (7K) 62.9
FT on Generated ($\sim$100K) 66.0
Self-consistency (Base, 32 paths) 77.1
At Least 1 Solve (Base, 32 paths) 95.6
Figure 1: An illustration of the reasoner-verifier paradigm. The verifier predicts scores for the solutions generated by the reasoner, and selectes the solution with the highest score.
Figure 2: We fine-tune PaLM 2-S on GSM8K. The base model is tuned with 2K training data to adhere with the problem format.
@@
Figure 3: The Model-induced Process Supervision (MiPS) data construction method we introduce in this work. By completing a intermediate solution with a reasoner several times, we can obtain the percentage value of these completions being correct, which is used to annotated the step-wise solutions.
@@
Training verifier models require solution wise or step wise labels, which is expensive to collect. There have been a series of work following an LLM-as-a-verifier approach where an off-the-shelf LLM is employed to judge the solutions through prompting Madaan et al. (2023); Kim et al. (2023); Pan et al. (2023). However, while such work may have seen improvements on language tasks, they haven’t been very successful in math or coding problems Huang et al. (2023); Luo et al. (2023b).
@@
In conclusion, to achieve optimal quality, training data is needed for building a strong verifier model. On the other hand, manually collecting solution verification labels is expensive and non-scalable. In this work, we propose to use Monte Carlo Sampling on the completions of the intermediate solutions to obtain step-wise training annotations (Fig. 3). Specifically, for each intermediate solution, we complete the solution with the reasoner several times through a sample decoding mechanism, and the percentage of the completed solutions being correct is referred to as the correctness of the solution. Such correctness scores can be used to train a PSV. Because of the nature of involving the reasoning model’s completion on the intermediate solutions, we call the construction of this data Model-induced Process Supervision (MiPS). While such an idea is also explored in a concurrent work Wang et al. (2023), we supplement with analysis of using MiPS constructed data. We found out that because the reasoner model, which completes the solutions, is not perfect, the noises it introduces would affect the design choices of using the process supervised verifier:
@@
  • We analyzed various ways to merge step-wise prediction scores to a single score value (we refer to this process as using an aggregation function) when using the verifier. Prior work used an aggregation function that focuses on low predicted scores that worked well for PSV trained on noise-free human annotated data Lightman et al. (2023). For the noisy MiPS data, we suggest aggregation functions that focuses on high predicted scores.
@@
  • We re-examine the usefulness of process supervision by isolating the trained PSV and studying the benefits of incorporating the score from each intermediate step during verification. Our results reveal that (1) the verification scores from later intermediate steps are indeed useful even for a PSV trained on the noisy MiPS data, however, the earlier step scores could hurt the verification, depending on the aggregation function; and (2) only using the PSV predicted score of the last step, in similar fashion as OSV, can sometimes be much better than OSV itself, indicating process supervision data can regularize OSV training.
@@
Following in this paper, we will provide a more complete review of related works, a precise definition of our method, and empirical results of the method and analysis on two math problem dataset and one coding dataset. The contributions of the paper is mainly (1) we propose MiPS to construct process supervision data automatically for training process supervision verifiers; (2) we extend the evaluation of problem solving verifiers to coding problems; (3) we provide empirical analysis on design choices and properties of the trained verifier from MiPS data.
@@ 2 Related Works
The advances of problem solving of LLMs can be broadly chracterized into two regimes, first by training a better reasoning model and the second by validating the solution from the reasoning model at inference time.
@@
Pre-training/Fine-tuning Better Reasoners. Standard training recipes also transfer to training better reasoners for problem solving. During pre-training, larger model sizes and training compute yields an LLM that is more competent in multiple aspects, including language abilities and problem solving (Achiam et al., 2023; Touvron et al., 2023; Anil et al., 2023, inter alia). Within fine-tuning, it is also observed that transfer learning Azerbayev et al. (2023) from a pile of generic math datasets, training on an augmented dataset on failure examples or diverse statements Huang et al. (2022); Luo et al. (2023a); Yu et al. (2023a); Ni et al. (2022) leads to improvements. Despite these approaches, it is apparent that (1) the state-of-the-art LLM still can fail at simple mistakes during multi-step problem solving and (2) the improvement of a simple verfication method by majority voting (self consistency Wang et al. (2022)) is still significant upon fine-tuning. Therefore, the exploration of verifiers to validate and pick the solutions is necessary.
@@
Validating Through LLM-as-a-verifier. There have been numerous attempts on using the LLM reasoner itself to correct and validate its generated solutions. Madaan et al. (2023); Kim et al. (2023) and many methods surveyed in Pan et al. (2023) broadly follow the strategy where the LLM validates and provides feedback to the generated solutions through prompting. Huang et al. (2023) and Luo et al. (2023b) revisited these methods and found that LLMs do not serve as good verifiers for equally competent solutions at the time being, and such methods improves marginally on math word problems.
@@
Validating Through Trained Verifiers. In contrast, verifiers trained on a human labelled dataset does show significant improvements Cobbe et al. (2021); Uesato et al. (2022); Lightman et al. (2023). Importantly, Lightman et al. (2023) showed that on a challenging competition-level mathematics problem set Hendrycks et al. (2021), verifiers trained on annotated intermediate solutions (PSV) surpasses verifiers trained on final solutions by a large margin, and both substantially better than self-consistency. Other analysis also emphasize on the importance of step-wise feedback: Uesato et al. (2022) showed that PSV selects solutions that are more accurate in their reasonings and Yao et al. (2023); Feng et al. (2023); Liu et al. (2023b), inter alia, showed that during decoding, LLMs can be guided towards better solutions step-by-step. We believe that improving the training of a PSV, and especially, identifying a scalable solution to generate the process supervision data, is of imminent importance. Therefore, in this work, we identify an automatic and generic solution to generate process supervision data (MiPS), and conducted detailed analysis centered on the noises of this automatic process.
@@
Math-Shepherd. Coincidentally, such an automatic process supervision data curation method was studied concurrently and independently by Wang et al. (2023). We share a generally similar methodology with their work, with a few minor design differences we highlight in later sections. The empirical results of MiPS is similar on the two datasets we share (GSM8K and MATH) despite using different, but about competent, LLMs. Their work extended training the verifier by applying it to fine-tune the reasoner through reinforcement learning, while our work included an additional coding dataset (MBPP) and provided analysis on the design choices of using the verifier, addressing the data noises. We believe these two works compliment each other.
@@ 3 Model-induced Process Supervision
We consider the reasoner-verifier framework where we start with a fairly competent reasoner on a task, generate verifier training data on a given set of problems with the reasoner, and train a verifier on the data to validate some new generated solutions by a reasoner. We first introduce discuss Model-induced Process Supervision, our data curation method that automatically creates process supervision data (annotations of accuracy of intermediate solutions). Then, we discuss the details about the verification process.
3.1 Scope
We start our introduction of MiPS with the requirements it is bound to: there should be a solution checker and the solutions could be decomposed into intermediate steps.
Without a solution checker, MiPS might require as much human annotation power as the original process supervision annotation schema Lightman et al. (2023). Fortunately, many tasks have a checker. For example, several math tasks are written in the form of multi-class classification (MMLU Hendrycks et al. (2020), MATHQA Amini et al. (2019)) or open-ended generation but with a special format that supports automatic parsing and checking (GSM8K Cobbe et al. (2021), MATH Hendrycks et al. (2021)). For coding problems, while the generated code is not directly verifiable, a common approach is to execute it on a series of test cases designed for the problem, and the code is only correct if it passes all the test cases (MBPP Austin et al. (2021), HumanEval Chen et al. (2021)).
Step-decomposition requires the solutions to be a multi-step problem that can be programmatically broken into intermediate solutions. This is also often the case. For example, in the math problems, the common approach is to treat a valid step as each individual line in the solution and the decomposition is simply dividing at the line breaks.
@@ 3.2 Obtaining MiPS data
MiPS constructs process supervision data through Monte Carlo sampling. First, we employ a reasoner model $r_{g}$ to generate a fix number of $n_{g}$ solutions for each problem, using temperature based decoding with a temperature of $t_{g}$. Then, for each solution, we decompose them into individual steps. After that, for each intermediate solution containing a prefix list of steps, we employ a reasoner model $r_{mc}$ to generate again a fix number of $n_{mc}$ solutions, with a termperature of $t_{mc}$, completing the intermediate solution. For each completed intermediate solution, we calculate the percentage (out of $n_{mc}$) of them being correct, and these correctness values comprises the MiPS data. In all experiments in this paper, we consider $r_{g}=r_{mc}$, namely, the reasoner model that is used to estimate the intermediate solution’s correctness is the same model that generates the solution data. This is particularly the most challenging case for MiPS, otherwise, using a more capable reasoner for the completion would enjoy a reduction of noise in MiPS data.
3.3 Training an Output Supervised Verifier (OSV)
We diverge a bit to illustrate how one would train a verifier from output supervision, which would later aid in explaining how the process supervised verifier is trained. The training data for OSV are the generations from the reasoner $r_{g}$ with the same temperature value $t_{g}$. The verifier itself is nothing different from a standard language model.
3.4 Training a Process Supervised Verifier (PSV)
The main differences of training PSV from OSV are:
  • To enable predicting a score at each step in the solution, we mark the last token of the each step (e.g., if each step is represented as a single line, the last token will be the new line token), and optimize step-wise predictions at each step at the same time. During inference, we would also obtain a score for each step in a solution.
  • While for the output supervision data, or human labelled process supervision data, the score is either 0 or 1, for MiPS data, the correctness scores are percentage values. The training objective considered in this work is to learn the percentage values $c_{i}$ for the $i$th step in the solution directly, through the same binary cross entropy loss. However, we note that it is possible to consider a different learning objective. For example, Wang et al. (2023) considered learning a binarized score:
@@ 3.5 Aggregating Step-wise Predictions
Table 1: We show the statistics of the datasets we use in this paper. The average number of steps is depicted with a granuarity of 0.5, using PaLM 2-S for GSM8K and MBPP, and PaLM 2-L for MATH. We note that these are not the most standard data splits, for reasons explained in Sec 4.2.
@@
They claimed that both are equivalently good aggregation functions. In later analysis, we show that for MiPS data, these two functions are underperforming for the trained verifier, whereas some tweaks
@@
performs much better. We analyze a much larger set of aggregation functions and suggest that the noise in the MiPS data causes the discrepancy.

### v2

@@ Abstract
Process supervision, using a trained verifier to evaluate the intermediate steps generated by a reasoner, has demonstrated significant improvements in multi-step problem solving. In this paper, to avoid the expensive effort of human annotation on the verifier training data, we introduce Model-induced Process Supervision (MiPS), a novel method for automating data curation. MiPS annotates an intermediate step by sampling completions of this solution through the reasoning model, and obtaining an accuracy defined as the proportion of correct completions. Inaccuracies of the reasoner would cause MiPS underestimating the accuracy of intermediate steps, therefore, we suggest and empirically show that verification focusing on high predicted scores of the verifier shall be preferred over that of low predicted scores, contrary to prior observations on human curated data. Our approach significantly improves the performance of PaLM 2 on math and coding tasks (accuracy +0.67% on GSM8K, +4.16% on MATH, +0.92% on MBPP compared with an output supervision trained verifier). Additionally, our study demonstrates that the verifier exhibits strong generalization ability across different reasoning models.
Multi-step Problem Solving Through a Verifier: An Empirical Analysis on Model-induced Process Supervision
Zihan Wang1thanks:   This work was done during the author’s internship at Google.      Yunxuan Li2thanks:   Correspondence: yunxuanli@google.com      Yuexin Wu2      Liangchen Luo2 Le Hou2   Hongkun Yu2      Jingbo Shang1 1University of California, San Diego        2Google
@@ 1 Introduction
Multi-step problem solving (e.g., math problems and coding challenges) showcases the capabilities of machine intelligence. While researchers have shown that model- and data-upscaling still hold powerful for large language models (LLMs) on multi-step problem solving Achiam et al. (2023); Touvron et al. (2023); Team Gemini et al. (2023); Huang et al. (2022); Azerbayev et al. (2023); Luo et al. (2023a); Yu et al. (2023b), even the state-of-the-art LLMs still produce easily observable mistakes. Furthermore, standard fine-tuning directly does not yield consistent and significant improvements Luo et al. (2023a); Yu et al. (2023b); Ni et al. (2022).
@@
Figure 1: An illustration of the reasoner-verifier paradigm. The verifier predicts scores for the solutions generated by the reasoner, and selects the solution with the highest score.
@@
Figure 2: The Model-induced Process Supervision (MiPS) data construction method we introduce in this work. By completing an intermediate solution with a reasoner several times, we can obtain the percentage value of these completions being correct. These annotations are used to train a process supervised verifier.
The reasoner-verifier paradigm (Fig. 1) is as an inference-time technique where the goal is to pick one model-generated solution among many, since it is observed that there often are some correctly generated solutions. In particular, self-consistency Wang et al. (2022) is a special case of the verifier that picks the solutions that shares the majority answer with others (e.g., math tasks where the answer is a number). LLM-based verifiers are more general, as they could apply to arbitrary text solutions (e.g., code that implements a function)
Training a verifier in a supervised fashion has demonstrated strong performance in both coding and math language problems. Cobbe et al. (2021) showed that by simply gathering correct and incorrect solutions to train a binary classification model and using such model to pick the highest confidence solution generated by the reasoner during inference time, the accuracy can be improved significantly. More recent studies suggest that verifying on intermediate steps could offer better guidance than training solely on the whole solutions Li et al. (2022); Uesato et al. (2022); Paul et al. (2023); Lightman et al. (2023); Feng et al. (2023); Yu et al. (2023a); Liu et al. (2023a); Wang et al. (2023). As such, verifiers trained (and applied) on intermediate steps are called process supervised verifiers (PSV), whereas those trained on whole solutions are called output supervised verifiers (OSV). In prior work, process supervision data is either obtained by an ad-hoc algorithm Li et al. (2022); Paul et al. (2023), or through expensive human annotations Uesato et al. (2022); Lightman et al. (2023), lacking an automatic and generic way of constructing of annotations of intermediate solutions.
@@
Training verifier models require solution wise or step-wise labels, which is expensive to collect. There have been a series of work following an LLM-as-a-verifier approach where an off-the-shelf LLM is employed to judge the solutions through prompting Madaan et al. (2023); Kim et al. (2023); Pan et al. (2023). However, while such work may have seen improvements on language tasks, they haven’t been very successful in math or coding problems Huang et al. (2023); Luo et al. (2023b).
@@
In conclusion, to achieve optimal quality, training data is needed for building a strong verifier model. On the other hand, manually collecting solution verification labels is expensive and non-scalable. In this work, we propose to use Monte Carlo Sampling on the completions of the intermediate solutions to obtain step-wise training annotations (Fig. 2). Specifically, for each intermediate solution, we complete the solution with the reasoner several times through a sample decoding mechanism, and the percentage of the completed solutions being correct is referred to as the correctness of the solution. The correctness scores are used to train a PSV. Because of the nature of involving the reasoning model’s completion on the intermediate solutions, we call the construction of this data Model-induced Process Supervision (MiPS). While such an idea is also explored in a concurrent work Wang et al. (2023), we supplement with analysis of using MiPS constructed data. We find that because the reasoner model, which completes the solutions, is not perfect, the noises it introduces would affect the design choices of training and using the process supervised verifier:
@@
  • We analyzed various ways to merge step-wise prediction scores to a single score value (we refer to this process as using an aggregation function) when using the verifier. Prior work used an aggregation function that focuses on low predicted scores and worked well for PSV trained on noise-free human annotated data Lightman et al. (2023). For the noisy MiPS data, we suggest aggregation functions that focus on high predicted scores.
@@
  • We re-examine the usefulness of process supervision by isolating the trained PSV and studying the benefits of incorporating the predicted score from each intermediate step during verification. Our results reveal that (1) the verification scores from later intermediate steps are indeed useful even for a PSV trained on the noisy MiPS data, however, the earlier step scores could hurt the verification; and (2) only using the PSV predicted score of the last step, in similar fashion as OSV, can sometimes be much better than OSV itself, indicating process supervision data can regularize OSV training.
@@
Following in this paper, we will provide a more complete review of related works, a precise definition of our method, and empirical results of the method and analysis on two math problem datasets and one coding dataset. The contributions of the paper are mainly (1) we propose MiPS to construct process supervision data automatically for training process supervision verifiers; (2) we extend the evaluation of problem solving verifiers to coding problems; (3) we provide empirical analysis on design choices and properties of the trained verifier from MiPS data.
@@ 2 Related Works
The advances of problem solving of LLMs can be broadly characterized into two regimes, first by training a better reasoning model and the second by validating the solution from the reasoning model at inference time.
@@
Pre-training/Fine-tuning Better Reasoners. Standard training recipes also transfer to training better reasoners for problem solving. During pre-training, larger model sizes and training compute yields an LLM that is more competent in multiple aspects (Achiam et al., 2023; Touvron et al., 2023; Anil et al., 2023, inter alia). Within fine-tuning, it is also observed that transfer learning Azerbayev et al. (2023) from a pile of generic math datasets, training on an augmented dataset of failure examples or diverse statements Huang et al. (2022); Luo et al. (2023a); Yu et al. (2023b); Ni et al. (2022); Khattab et al. (2023) leads to improvements. Despite these approaches, it is apparent that (1) the state-of-the-art LLM still can fail at simple mistakes during multi-step problem solving and (2) the improvement of a simple verification method by majority voting (self-consistency Wang et al. (2022)) is still significant upon fine-tuning. Therefore, the exploration of verifiers to validate and pick the solutions is necessary.
@@
Validating Through LLM-as-a-verifier. There have been numerous attempts on using the LLM reasoner itself to correct and validate its generated solutions. Madaan et al. (2023); Kim et al. (2023) and many methods surveyed in Pan et al. (2023) broadly follow the strategy where the LLM validates and provides feedback to the generated solutions through prompts. Huang et al. (2023) and Luo et al. (2023b) revisited these methods and found that LLMs are not good verifiers for equally competent solutions, as such methods improves marginally on math word problems.
@@
Validating Through Trained Verifiers. In contrast, verifiers trained on a human labelled dataset does show significant improvements Cobbe et al. (2021); Uesato et al. (2022); Lightman et al. (2023). Importantly, Lightman et al. (2023) showed that on a challenging competition-level mathematics problem set Hendrycks et al. (2021), verifiers trained on annotated intermediate solutions (PSV) surpasses verifiers trained on final solutions by a large margin, and both substantially better than self-consistency. Other analysis also emphasized on the importance of step-wise feedback: Uesato et al. (2022) showed that PSV selects solutions that are more accurate in their reasoning and Yao et al. (2023); Feng et al. (2023); Liu et al. (2023b), inter alia, showed that during decoding, LLMs can be guided towards better solutions step-by-step. We believe that improving the training of a PSV, and especially, identifying a scalable solution to generate the process supervision data, is of imminent importance. Therefore, in this work, we identify an automatic and generic solution to generate process supervision data (MiPS), and conducted detailed analysis centered on the noises of this automatic process.
@@
Math-Shepherd. Coincidentally, such an automatic process supervision data curation method was studied concurrently and independently by Wang et al. (2023). We share a generally similar methodology with their work, with a few minor design differences we highlight in later sections. The empirical results of MiPS are similar on the two datasets we share (GSM8K and MATH) despite using different, but about competent, LLMs. Their work extended training the verifier by applying it to fine-tune the reasoner through reinforcement learning, while our work included an additional coding dataset (MBPP) and provided analysis on the design choices of using the verifier, addressing the data noises. We believe these two works complement each other.
@@ 3 Model-induced Process Supervision
We consider the reasoner-verifier framework where we start with a fairly competent reasoner on a task, generate verifier training data on a given set of problems with the reasoner, and train a verifier on the data to validate some new generated solutions by a reasoner. We first discuss Model-induced Process Supervision, our data curation method that automatically creates process supervision data. Then, we discuss the details about the verification process.
@@ 3.1 Obtaining MiPS data
MiPS constructs process supervision data through Monte Carlo sampling. First, we employ a reasoner model $r_{g}$ to generate a fix number of $n_{g}$ solutions for each problem, using temperature based decoding with a temperature of $t_{g}$. Then, for each solution, we decompose them into individual steps (we treat each line in a solution as an individual step). After that, for each intermediate solution containing a prefix list of steps, we employ a reasoner model $r_{mc}$ to generate again $n_{mc}$ solutions, with a temperature of $t_{mc}$, completing the intermediate solution. For each completed intermediate solution, we calculate the percentage (out of $n_{mc}$) of them being correct, and these correctness values comprises the MiPS data. In all experiments in this paper, we consider $r_{g}=r_{mc}$, namely, the reasoner model that is used to estimate the intermediate solution’s correctness is the same model that generates the solution data. This is particularly the most challenging case for MiPS, otherwise, using a more capable reasoner for the completion can enjoy a reduction of noise in MiPS data.
3.2 Training an Output Supervised Verifier
To understand how well the process supervised verifier (PSV) trained from MiPS is, it is necessary to consider the vanilla output supervised verifier (OSV), which uses the same amount of human labeling resources. The training data for OSV are the generations from the reasoner $r_{g}$ with the same temperature value $t_{g}$. The verifier itself is a standard language model that supports a binary classification on the final token of the input. The verifier is trained on the cross entropy loss of the prediction. This is also known as the solution-level verifier in Cobbe et al. (2021).
3.3 Training a Process Supervised Verifier
The differences of training PSV and OSV are:
  • To enable predicting a score at each step in the solution, we mark the last token of each step (e.g., if each step is represented as a single line, the last token will be the new line token), and optimize step-wise predictions at each step at the same time. During inference, we would also obtain a score for each step in a solution.
  • While for the output supervision data, or human labelled process supervision data, the score is either 0 or 1, for MiPS data, the correctness scores are percentage values. The training objective considered in this work is to learn the exact percentage values $c_{i}$ for the $i$th step in the solution directly. However, we note that it is possible to consider a different learning objective. For example, Wang et al. (2023) considered learning a binarized score:
@@ 3.4 Aggregating Step-wise Predictions
Table 1: We show the statistics of the datasets we use in this paper. The average number of steps is depicted with a granularity of 0.5, using PaLM 2-S for GSM8K and MBPP, and PaLM 2-L for MATH. We note that these are not the most standard data splits, for reasons explained in Sec 4.2.
@@
They claimed that both are equivalently good aggregation functions. In later analysis, we show that for MiPS data, these two functions are underperforming for the trained verifier, while,
@@
are much better. We provide an analysis with a much larger set of aggregation functions and suggest that MiPS data prefers aggregation functions that focus on high prediction scores rather than lower ones.